OpenStudy (anonymous):

I have the following diff eq: xdy + (xy+y-1)dx =0 My attempt to solve: Since they are not exact I tried to find the integrating factor by using exp{S 1/M [ dN/dx - dM/dy)dy to which I got e^x. I am supposed to multiply this integrating factor with the original eq. but this doesn't take me anywhere. I am stuck.. please help! Thanks.

OpenStudy (anonymous):

xdy+xydx+ydx-dx=0 xydx+ydx-dx=-xdy dx(xy+y-1)=-xdy xy+y-1=-x(dy/dx) (xy+y-1)/(-x)=dy/dx -y-(y/x)+(1/x)=dy/dx

OpenStudy (anonymous):

or if you only wanted y' xy'+xyx'+yx'-x'=0 y'=(xyx'+yx'-x')/-x y'=-yx'-(y/x)x'+(1/x)x'

OpenStudy (anonymous):

well I was supposed to use an integrating factor and the convert it into an exact equation... but thanks for the help